ICU Window, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 10 years to maturity that is quoted at 104.5 percent of face value. The issue makes semiannual payments and has an embedded cost of 5.6 percent annually.

a. What is the company’s pretax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b. If the tax rate is 23 percent, what is the aftertax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a. Semi annual coupon =1000*5.6%/2 =56/2=28
Number of semi annual periods =10*2 =20
Price =104.5%*1000 =1045
Par value =1000
Pretax cost of debt =2*RATE(20,28,-1045,1000) =5.0221%

b. After tax cost of debt =5.0221%*(1-23%) =3.87%